4. A park has a stable population of birds. Prior to this situation, the birds’ population

increased from an initial low level. When the population of birds was 1000, the

proportionate birth rate was 40% per year and the proportionate death rate was 5% per

year. When the population was 3,000, the proportionate birth rate was 30% and the

proportionate death rate was 10%. Consider the population model under the following

assumptions: (10)

(i) There is no migration and no exploitation.

(ii) The proportionate birth rate is a decreasing linear function of the population.

(iii) The proportionate death rate is an increasing linear function of the population.

Show that

The population grows according to the logistic model.

Find the stable population size.

If the shooting of birds is allowed at the rate of 15% of the population per year, find the

new equilibrium population.


1
Expert's answer
2021-04-27T13:01:32-0400


1)"\frac{dN}{dt}=(\beta-\delta)N"


"N(t)-population \space size"

"\beta(N)-" birth rate

"\delta(N)-" death rate


"N=1000\implies \beta =40"

"N=3000\implies \beta =30"



here "\beta" is birth rate



here we have seen 1000 is population when birth rate is 40%

and also

here we have seen 3000 is population when birth rate is 5%

so equation comes out to be

"\beta(N)=-0.005N+45"




"N=1000\implies \delta=5"

"N=3000\implies \delta=10"



here "\delta" is death rate

here we have seen 1000 is population when death rate is 5%

and also

here we have seen 3000 is population when death rate is 10%

"\delta(N)=-0.0025N+2.5"



"\beta-\delta =
\frac{dN}{dt}"


"\beta-\delta = -0.0075N-42.5" ......................(a)

putting (a) = 0 to get equilibrium points

i.e

"\frac{dN}{dt}=0"

"(-0.0075N + 42.5)N=0"

Equilibrium points :-

N= 0 , N≈3667........(1)




"2) \n\n{dN\\over dt}\n\u200b\t\n =(\u22120.0075N+42.5\u221215)N"


"{\ndN\\over dt}\n\u200b\t\n =(\u22120.0075N+27.5)N" this is the equation that will come when we subtract the shooting rate of 15 , this is new equilibrium population.



Equilibrium points :- from (1) for part 1 answer

N=0, N≈3667